OpenStudy (anonymous):

Find a six-digit multiple of 64 that consists only of the digits 1 and 2.

OpenStudy (anonymous):

If it helps any 100032 is the smallest 6 digit multiple of 64

OpenStudy (anonymous):

Oh. and 111104 is the smallest multiple where there are no more zeroes in the first 4 digits.

OpenStudy (anonymous):

Just sharing information. The answer is going to be between \(64\cdot 1563\) and \(64\cdot 4687\). After that you get into the 300 thousands, which wont satisfy the conditions.

OpenStudy (anonymous):

The ones digit must also be a 3 or an 8.

OpenStudy (anonymous):

But, it has to consist of only 1 and 2

OpenStudy (anonymous):

Er, the ones digit of the number that you multiply with 64.

OpenStudy (cwrw238):

i guess if you keep adding 320 to 100032 you\ll eventually get there

OpenStudy (anonymous):

That's what I meantm my bad >.<

OpenStudy (anonymous):

ooh.

OpenStudy (anonymous):

cwrw is correct.

OpenStudy (cwrw238):

a bit tedious that...

OpenStudy (anonymous):

it is isn't it? There has to be a simpler way...

OpenStudy (anonymous):

Yeah, I want to believe there is a more though oriented way, instead of guess and check.

OpenStudy (anonymous):

I got the answer btw <.< totally guess and check though.

OpenStudy (anonymous):

The multiple is 122112, its \(64\cdot 1908\)

OpenStudy (loser66):

345658

OpenStudy (loser66):

345658*64= 22122112

OpenStudy (anonymous):

I just realized, because there can only be 1's and 2's in the result, and its 6 digits long, there are only \(2^6\) possible candidates.

OpenStudy (anonymous):

Take away the fact that 1 can't be the units digit, now there are \(2^5\) possible candidates.

OpenStudy (slaaibak):

Well, firstly, if a number is divisible by 64, it is divisible by 8. If a number is divisible by 8, the last three digits should be divisible by 8. Firstly, it must end with 2. So the number is in the form ......2 so the last three digits is either 112, 122, 212, 222 Only 112 is divisible by 8. So now we know the number is of the form ...112 Moreover, the number should also be divisible by 16. Now the last 4 digits have to be divisible by 16. So it can either be ..2112 or ..1121 Only 2112 is divisible by 16. So our number is ..2112 Now just test the last 4 possibilities (either 112112, 122112, 212112, 222112)

OpenStudy (anonymous):

Very interesteing. :) You could keep going with 32 and 64, then you would be done.

OpenStudy (slaaibak):

Sorry, the fourth last line should read: So it can either be ..2112 or ..1112

OpenStudy (cwrw238):

great stuff

OpenStudy (anonymous):

112112 doesnt work

OpenStudy (slaaibak):

122112 does work though.

OpenStudy (anonymous):

its 122112. Because 22112 is divisible by 32, but 12112 isnt.

OpenStudy (anonymous):

it does! yay!

OpenStudy (anonymous):

Thanks everyone!

OpenStudy (anonymous):

I wish i could give medals to alll of you XD

OpenStudy (slaaibak):

A reason why a 4 or higher digit number is divisible by 8 iff the last 3 digits is divisible by 8: any four digit number (or higher) can be written as axyz Now, axyz can be written as 1000a + xyz We know 1000 is divisible by 8. So if axyz = 1000a + xyz is divisible by 8, it implies xyz is divisible by 8. This follows, because if any terms which are divisible by a number are added, the sum is also divisible by that number.